The work done by an external force to move a particle of charge -2.20 uC at constant speed from point A to point B is 6.40x10-3J. (a) What is the difference between the electric potential energy of the charge at the two points? (b) What is the difference in the kinetic energy of the charge at the two points? (c) Determine the potential difference between the two points. (d) State which point is at the higher potential.
Answer #1

a)

From work energy

work done = change in potential energy

therefore,

the difference between the electric potential energy of charge at two points =6.40*10^-3 J

b)

delta K = Kf - Ki

but Vf=Vi

so, delta K = 0 J

c)

delta U =q*delta V

delta V= delta U/(q)

= 6.40*10^-3 J/(2.20*10^-6 C)

= 2909.090909 volts

delta V = 2909 volts

d)

point A is at higher potential than B.Since it requires positive work to move the negative charge from A to B.